If volume of given mass of a gas is doubled at atmospheric pressure then temperature of gas rises from 127 degree C to ? a.527degC b.572degC c.500degC d. 625degC

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
As V proportional to T
Vf = 2 * Vi
Vf / Tf = Vi / Ti
2 Vi / Tf = Vi / 400
Tf = 800 K
Tf = 527 degree Centigrade